OpenStudy (anonymous):

Let \(E\subseteq\ \mathbb{R}\) Prove that: a) \(E\) is open iff \(E^{o} = E\) b) If \(O\subseteq\ E\) and \(O\) is open, then \(O \subseteq\ E^{o}\) c) If \(E\) is dense in \(\mathbb{R}\), then \(\forall\ p\ \in\ \mathbb{R},\ \exists\) a sequence \(p_{n} \in\ E\) s.t \(\lim_{n \rightarrow \infty} p_{n} = p\)

OpenStudy (anonymous):

\(E^{o}\) is the set of all interior points of E. No idea if that's the conventional notation or not.

OpenStudy (zzr0ck3r):

What is your definition of open?

OpenStudy (zzr0ck3r):

is the epsilon ball definition?

OpenStudy (anonymous):

Hmm, Rudin?

OpenStudy (anonymous):

Um " A subset of \(E\) of \(\mathbb{R}\) is said to be open if \(E = E^{o}\) Not Rudin, it's something my professor made on his own. He combined Rudin and two other textbooks and picked bits and pieces out of each, apparently.

OpenStudy (anonymous):

Any luck on any of these proofs?

OpenStudy (anonymous):

I havent looked at them as much as I would like. So I'm still in the brainstorming stage :P

OpenStudy (anonymous):

I guess the first one seems odd to me because it seems just be restating the definition I have given to me. By definition, E is open if \(E = E^{o}\). So to ask me to prove that seems weird because it seems to be true simply by definition, not by using theorems and other knowledge I have access to.

OpenStudy (anonymous):

If that's the case, you have half the proof given to you.

OpenStudy (anonymous):

I suppose the other half would require me to show the converse?

OpenStudy (anonymous):

I should've caught on to that, the definition saying if, the proof asking iff, lol.

OpenStudy (anonymous):

Right. The way I had it introduced to me is that an open set is a set whose points are all interior points. This means if \(E\) is open, then \(E\subset E^o\). Since a set's interior points belong to the set, you also have \(E^o\subset E\), and so \(E=E^o\).

OpenStudy (anonymous):

I guess conceptually then, that's only possible for an infinite set then, right?

OpenStudy (anonymous):

Not necessarily, according to this link: http://mathworld.wolfram.com/OpenSet.html "... it is possible for a more general topological set to be both finite and open."

OpenStudy (anonymous):

Alright, so Ill make sure to keep that in mind before I just throw out that claim. Hm...I think I have an idea for the 2nd one, but it kind of revolves around the idea that if a set is infinite, then it is open. But again, not sure I can just claim that. If I can then Im sure I have #2.

OpenStudy (anonymous):

I suppose its similar to my previous question. Just instead of saying open implies infinite, I guess im hopign I can claim the opposite, infinite implies open.

OpenStudy (anonymous):

Do you have the converse of the first?

OpenStudy (anonymous):

What do you mean? As in do I know how to phrase it?

OpenStudy (anonymous):

How to prove it. It's fairly simple, straight from the definition of an open set.

OpenStudy (anonymous):

Well, if O is open, then it contains all of its interior points. And since O is a subset of E, all of the interior points of O are contained in E. But then the next step, I think, would require that E is always a subset of \(E^{o}\) which I guess we just said above. But....it doesnt explicitly say it in my notes, but I guess that has to be true, eh? \(E \subseteq\ E^{o}\) is always true?

OpenStudy (anonymous):

Like, I guess you can't ever say \(|E^{o}| < |E|\)?

OpenStudy (anonymous):

Hmm, I think you're going about it wrong. Try showing that \(O^o\subset E^o\). Then you'll know that since \(O\) is open, you have \(O=O^o\) by the first proof, and so \(O\subseteq E^o\).

OpenStudy (anonymous):

Right. I guess I thought that required me to be certain that \(E \subseteq\ E^{o}\) was always true. I wasn't sure I could make the jump from \(O \subseteq\ E\) to \(O^{o} \subseteq\ E^{o}\) Without that n between step. Um....Oh, but you dont want me to go from \(O\) to \(O^{o}\), you think I need to start with \(O^{o}\)?

OpenStudy (anonymous):

Start off with some interior point \(p\) of \(O\). Using the definition of an interior point, you have a neighborhood of \(p\), call it \(N\), where \(N\subset O\). Now since \(O\subset E\), you have \(N\subset E\), so \(p\in E^o\). This means \(O^o\subset E^o\). The rest of the proof follows from here.

OpenStudy (anonymous):

Okay, using the definition of interior point let's me use the neighborhood idea, which will allow me to safely say my chosen element is in \(E^{o}\) I think I feel more safely about that then. That probably outlines that well enough. I guess now I'm trying to understand the idea of "dense." The way it's defined for me is: "Let E be a subset of \(\mathbb{R}\). E is said to be dense in \(\mathbb{R}\) if \(E \cup\ E' = \mathbb{R}\)." So, since E' is the set of all limit points of E, I would think that all \(E \cup\ E'\) really means is that you must include the endpoints of the set E. Which I guess reads to be as in you need a closed interval that equals \(\mathbb{R}\). But I know this interpretation is wrong since it is noted that \(\mathbb{Q}\) is dense in R. So maybe I'm misunderstanding limit points or something, but I'm missing something about the definition.

OpenStudy (anonymous):

I never really had a good idea of what dense meant myself, or maybe I did but it's been too long... in any case, I suggest watching these lectures, they really helped me out: http://www.youtube.com/watch?v=Ebnoxgp8mLM http://www.youtube.com/watch?v=6zs_PTUfKAk The prof doesn't really delve into the definition of dense until the last 4 or 5 minutes of the second lecture, but it might give you a better idea of what dense means. (The other parts of the videos make up a really nice comprehensive review of the other definitions.)

OpenStudy (anonymous):

Alright, Ill look them over and go from there. Thanks as always, Sith :)

OpenStudy (anonymous):

Happy to help!

OpenStudy (anonymous):

^_^